# On the Performance Analysis of Momentum Method: A Frequency Domain Perspective

### v1

@@ B.3 Additional Experiments of VGG16 on CIFAR-10
B.4 The Early Stages of Training
@@
Fig. 8 demonstrates the early test accuracy curves of different momentum coefficient methods. For orthodox momentum systems, preserving the original gradient (i.e., all-pass momentum system, low-pass momentum system with an increasing $u_{t}$, and high-pass momentum system with an increasing $u_{t}$) or attenuating high-frequency gradient components(i.e., static low-pass momentum system with $u_{t}=0.9$) results in better initial performance, while greatly attenuating high-frequency gradient components (i.e., low-pass momentum system with a decreasing $u_{t}$) or attenuating low-pass components (i.e., static high-pass and high-pass momentum system with a decreasing $u_{t}$) lead to bad test performance at the beginning.
@@
B.5 Comparison with Special Momentum Systems
@@
In this subsection, we investigate the test performance of the following four types of momentum systems: 1) low-pass to high-pass momentum system (LP2HP); 2) high-pass to low-pass momentum system (HP2LP); 3) low-pass gain to high-pass gain momentum system (LPG2HPG); 4) high-pass gain to low-pass gain momentum system (HPG2LPG). Their dynamic magnitude responses are shown in Fig. 9. Note that the maximum values $|H(\omega)|$ of these four systems are the same as the default setting in FSGDM. We run each experiment under 3 different random seeds (0-2). Table 8 displays the test accuracy results of four types of momentum systems and FSGDM. Clearly, our proposed FSGDM outperforms all four special momentum systems. Specifically, the test accuracy of the momentum systems shifting from high-pass to low-pass is better than that shifting from low-pass to high-pass. This indicates that compared to the low-frequency gradient components, high-frequency components are more undesired in the late training stages, which supports the finding in Section 3.
@@
B.6 Training with Extreme Momentum Coefficients
@@
Why do researchers usually choose $u_{t}=0.9$ or $v_{t}=1$ instead of larger values? From the frequency domain perspective, we discover that 1) when $u_{t}$ is extremely close to $1$ in EMA-SGDM, the momentum system will behave like a super narrow low-pass filter, with an extreme reduction in most of the high-frequency gradient components; 2) when $u_{t}$ is extremely close to $1$ in Standard-SGDM, the momentum system will behave like a super narrow low-pass gain filter, with a reduction in high-frequency gradient components and high amplification in a narrow band of low-frequency gradient components; 3) when $v_{t}$ is larger than $1$ in Standard-SGDM, the attenuation of high-frequency gradient components is then reduced. We speculate that all these poor filtering characteristics of the momentum systems will lead to bad test performance. Fig. 10 displays the magnitude response of these three situations. As shown in Fig. 11, the test performance results validate our previous speculations and support our frequency domain analysis framework.
@@
B.7 Additional Exploration of Optimal Settings For NLP Tasks
In this subsection, we provide experiments that explore the optimal parameter selection of FSGDM for the IWSLT14 translation task by training LSTM-W and Transformer-tiny. The experimental settings follow Section 5 and Appendix C.
@@ B.7 Additional Exploration of Optimal Settings For NLP Tasks
The results in Fig. 12 indicate that similar optimal zones can be observed on the NLP task. When the momentum coefficient $v$ is fixed, the BLEU score shows an initial increase followed by a decline as the scaling factor $c$ increases, which is highly consistent with the results in Section 4.2. In addition, we find that the empirical insights discussed in Section 3.3 are also applicable to various deep learning models beyond CNNs, as well as NLP tasks.
B.8 Optimal Zone of FSGDM
@@
In this subsection, we delve further into the optimal zone. We suspect that the similarity of the dynamic magnitude responses may lead to close test set performance. The dynamic magnitude responses of the black points with different parameters in the optimal zone (Fig. 4) are shown in Fig. 13. We train ResNet50 on CIFAR-100 and visualize the training losses and the test accuracy curves of different points in the optimal zone. The results are shown in Fig. 14.
@@
B.9 Ablation Study on Different Batch Size
This subsection provides the experiments of training ResNet50 on CIFAR-100 with different batch size settings. We compare the Top-1 accuracy of the test set by using our FSGDM with $c=0.033,v=1$, Standard-SGDM with $u_{t}=0.9,v_{t}=1$, and EMA-SGDM with $u_{t}=0.9$, as shown in Table 9. The test results show that our FSGDM consistently outperforms popular conventional SGD-based momentum optimizers.
@@ B.9 Ablation Study on Different Batch Size
Appendix C Experimental Settings
C.1 Training Settings for Vision Classification Tasks
We use custom training code based on the PyTorch tutorial code for all our visual classification experiments (including the experiments in Section 3, Section 4.2 and Section 5) We choose the CosineAnnealingLR (Loshchilov and Hutter,, 2016) as our training scheduler. Additionally, we set the learning rate as $1\times 10^{-1}$ for all experiments, while the weight decay is set as $5\times 10^{-4}$ for experiments on CIFAR-10, CIFAR-100 and Tiny-ImageNet, and $1\times 10^{-1}$ for ImageNet. All models we used are simply following their paper’s original architecture, and adopt the weight initialization introduced by He et al., (2015). Additionally, we train 300 epochs for experiments on CIFAR-10 and CIFAR-100 and train 100 epochs for Tiny-ImageNet and ImageNet. We use a 128 batch size for experiments on CIFAR-10, CIFAR-100 and Tiny-ImageNet, and 256 for ImageNet. All experiments are conducted on RTX 4090 or A100 GPUs.
Data Augmentation. For experiments on CIFAR-10, CIFAR-100 and Tiny-ImageNet, we adopt PyTorch’s RandomCrop, followed by random horizontal flips. Specifically, the random crop size is set to 32x32 for CIFAR-10 and CIFAR-100 and set to 64x64 for Tiny-ImageNet. For experiments on ImageNet, we adopt PyTorch’s RandomResizedCrop, cropping to 224x224 followed by random horizontal flips. Test images use a fixed resize to 256x256 followed by a center crop to 224x224. At last, a data normalization is adopted to input images.
C.2 Training Settings for Natural Language Processing Tasks
@@
C.3 Training Settings for Reinforcement Learning Tasks
For the experiments in RL tasks, we do not make any changes except for replacing the original Adam optimizer with Standard-SGDM, EMA-SGDM and our proposed FSGDM. To ensure fairness, we use Tianshou’s (Weng et al.,, 2022) default hyperparameters for PPO training. However, since SGD-based optimizers are highly sensitive to the learning rate, we conducted a search for suitable learning rates across the three games, ultimately setting $10^{-2}$, $10^{-2}$ and $10^{-3}$ for Walker2d-v4, HalfCheetah-v4, and Ant-v4, respectively.
Appendix D Challenges in the Frequency Domain Analysis for Adaptive Optimizers
Input $\beta_{2}$, $\epsilon$;
Initialization $v_{0}$;
@@ Appendix D Challenges in the Frequency Domain Analysis for Adaptive Optimizers
Input $\beta_{1}$, $\beta_{2}$, $\epsilon$;
Initialization $m_{0}$, $v_{0}$;
@@
At this stage, we suspect that our argument regarding the three insights discussed in Section 3.3 might be also applicable to other types of optimizers. However, how the different frequency gradient components in the model parameter updates are actually processed by the Adam optimizer remains unclear. We anticipate that resolving these issues will provide deeper insights.

### v3

@@ C.3 Additional Experiments of VGG16 on CIFAR-10
C.4 The Early Stages of Training
@@
Figure 8 demonstrates the early test accuracy curves of different momentum coefficient methods. For orthodox momentum systems, preserving the original gradient (i.e., all-pass momentum system, low-pass momentum system with an increasing $u_{t}$, and high-pass momentum system with an increasing $u_{t}$) or attenuating high-frequency gradient components(i.e., static low-pass momentum system with $u_{t}=0.9$) results in better initial performance, while greatly attenuating high-frequency gradient components (i.e., low-pass momentum system with a decreasing $u_{t}$) or attenuating low-pass components (i.e., static high-pass and high-pass momentum system with a decreasing $u_{t}$) lead to bad test performance at the beginning.
@@
C.5 Comparison with Special Momentum Systems
@@
In this subsection, we investigate the test performance of the following four types of momentum systems: 1) low-pass to high-pass momentum system (LP2HP); 2) high-pass to low-pass momentum system (HP2LP); 3) low-pass gain to high-pass gain momentum system (LPG2HPG); 4) high-pass gain to low-pass gain momentum system (HPG2LPG). Their dynamic magnitude responses are shown in Figure 9. Note that the maximum values $|H(\omega)|$ of these four systems are the same as the default setting in FSGDM. We run each experiment under 3 different random seeds (0-2). Table 8 displays the test accuracy results of four types of momentum systems and FSGDM. Our proposed FSGDM outperforms all four special momentum systems. Specifically, the test accuracy of the momentum systems shifting from high-pass to low-pass is better than that shifting from low-pass to high-pass. This indicates that compared to the low-frequency gradient components, high-frequency components are more undesired in the late training stages, which supports the finding in Section 3.
@@
C.6 Training with Extreme Momentum Coefficients
@@
Why do researchers usually choose $u_{t}=0.9$ or $v_{t}=1$ instead of larger values? From the frequency domain perspective, we discover that 1) when $u_{t}$ is extremely close to $1$ in EMA-SGDM, the momentum system will behave like a super narrow low-pass filter, with an extreme reduction in most of the high-frequency gradient components; 2) when $u_{t}$ is extremely close to $1$ in Standard-SGDM, the momentum system will behave like a super narrow low-pass gain filter, with a reduction in high-frequency gradient components and high amplification in a narrow band of low-frequency gradient components; 3) when $v_{t}$ is larger than $1$ in Standard-SGDM, the attenuation of high-frequency gradient components is then reduced. We speculate that all these poor filtering characteristics of the momentum systems will lead to bad test performance. Figure 10 displays the magnitude response of these three situations. As shown in Figure 11, the test performance results validate our previous speculations and support our frequency domain analysis framework.
@@
C.7 Additional Exploration of Optimal Settings For NLP Tasks
In this subsection, we provide experiments that explore the optimal parameter selection of FSGDM for the IWSLT14 translation task by training LSTM-W and Transformer-tiny. The experimental settings follow Section 5 and Appendix D.
@@ C.7 Additional Exploration of Optimal Settings For NLP Tasks
The results in Figure 12 indicate that similar optimal zones can be observed on the NLP task. When the momentum coefficient $v$ is fixed, the BLEU score shows an initial increase followed by a decline as the scaling factor $c$ increases, which is highly consistent with the results in Section 4.2. In addition, we find that the empirical insights discussed in Section 3.3 are also applicable to various deep learning models beyond CNNs, as well as NLP tasks.
C.8 Optimal Zone of FSGDM
@@
In this subsection, we go deeper into the optimal zone. We suspect that the similarity of the dynamic magnitude responses may lead to close test set performance. The dynamic magnitude responses of the black points with different parameters in the optimal zone (Figure 4) are shown in Figure 13. We train ResNet50 on CIFAR-100 and visualize the training losses and the test accuracy curves of different points in the optimal zone. The results are shown in Figure 14.
@@
C.9 Ablation Study on Different Batch Size
This subsection provides the ResNet50 training experiments on CIFAR-100 with different batch size settings. We compare the Top-1 accuracy of the test set by using our FSGDM with $c=0.033,v=1$, Standard-SGDM with $u_{t}=0.9,v_{t}=1$, and EMA-SGDM with $u_{t}=0.9$, as shown in Table 9. The test results show that our FSGDM consistently outperforms popular conventional SGD-based momentum optimizers.
@@ C.9 Ablation Study on Different Batch Size
Appendix D Experimental Settings
D.1 Training Settings for Vision Classification Tasks
We use custom training code based on the PyTorch tutorial code for all our visual classification experiments (including the experiments in Section 3, Section 4.2 and Section 5) We choose the CosineAnnealingLR (Loshchilov & Hutter, 2016) as our training scheduler. Additionally, we set the learning rate as $1\times 10^{-1}$ for all experiments, while the weight decay is set as $5\times 10^{-4}$ for experiments on CIFAR-10, CIFAR-100, and Tiny-ImageNet, and $1\times 10^{-1}$ for ImageNet. All models we used are simply following their paper’s original architecture, and adopt the weight initialization introduced by He et al. (2015). Additionally, we train 300 epochs for experiments on CIFAR-10 and CIFAR-100 and train 100 epochs for Tiny-ImageNet and ImageNet. We use a 128 batch size for experiments on CIFAR-10, CIFAR-100, and Tiny-ImageNet, and 256 for ImageNet. All experiments are conducted on RTX 4090 or A100 GPUs.
Data Augmentation. For experiments on CIFAR-10, CIFAR-100, and Tiny-ImageNet, we adopt PyTorch’s RandomCrop, followed by random horizontal flips. Specifically, the random crop size is set to 32x32 for CIFAR-10 and CIFAR-100 and set to 64x64 for Tiny-ImageNet. For experiments on ImageNet, we adopt PyTorch’s RandomResizedCrop, cropping to 224x224 followed by random horizontal flips. Test images use a fixed resize to 256x256 followed by a center crop to 224x224. At last, a data normalization is adopted to input images.
D.2 Training Settings for Natural Language Processing Tasks
@@
D.3 Training Settings for Reinforcement Learning Tasks
For the experiments in RL tasks, we do not make any changes except for replacing the original Adam optimizer with Standard-SGDM, EMA-SGDM, and our proposed FSGDM. To ensure fairness, we use Tianshou’s (Weng et al., 2022) default hyperparameters for PPO training. However, since SGD-based optimizers are highly sensitive to the learning rate, we searched for suitable learning rates across the three games, ultimately setting $10^{-2}$, $10^{-2}$ and $10^{-3}$ for Walker2d-v4, HalfCheetah-v4, and Ant-v4, respectively.
Appendix E Challenges in the Frequency Domain Analysis for Adaptive Optimizers
Input $\beta_{2}$, $\epsilon$, $v_{0}$;
@@ Appendix E Challenges in the Frequency Domain Analysis for Adaptive Optimizers
Input $\beta_{1}$, $\beta_{2}$, $\epsilon$, $m_{0}$,$v_{0}$;
@@
At this stage, we believe that our argument regarding the three insights discussed in Section 3.3 is also applicable to other optimizers. However, it remains unclear how the different frequency gradient components in the model parameter updates are processed by the Adam optimizer. We anticipate that resolving these issues will provide deeper insight.